Derrick is 4 times older than Sarah 3 years ago the sum of their age was 24 how old are each now?
Answer by Cromlix(4381)

Derrick = 4 * Sarah (4S)
D - 3 + S - 3 = 24
Replace D with 4S
4S - 3 + S - 3 = 24
Collect like terms:
5S - 6 = 24
5S = 24 + 6
5S = 30
S = 6
Sarah is 6 years old.
Derrick is 24 years old.
Hope this helps.
:-)